# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 3.2.1 Bohr’s Quadrant: Nash Implementation
Maskin examines the following abstract setting: There is a set of individuals $I$ and a set of outcomes $A$, where each individual $i\in I$ has a weak preference relation $R_{i}$ over $A$. For each individual $i\in I$, let $\mathcal{R}_{i}$ represent the set of possible preferences. Let $R=(R_{i})_{i\in I}$ denote a generic preference profile, and let $\mathcal{R}$ denote the set of all possible preference profiles. Importantly, preferences are private information.
A central planner aims to select an outcome based on individuals’ true preferences, formalized as a social choice function $f:\mathcal{R}\to A$ that maps each preference profile $R\in\mathcal{R}$ to an outcome $\alpha\in A$ [^16] Maskin’s theory is more general, allowing for social choice correspondences that map each preference profile to a set of outcomes.
@@
To overcome this, the planner designs a mechanism $(S,g)$ la Hurwicz, where $S_{i}$ represents the strategy space for each individual $i\in I$, and $g:S\to A$ is the outcome function selecting an outcome $\alpha\in A$ for each strategy profile $s=(s_{i})_{i\in I}\in S=\prod_{i\in I}S_{i}$. The objective is to design a mechanism such that, for any preference profile $R$, the Nash equilibrium outcomes of the induced game coincide with the desired outcome $f(R)$. This allows the planner to implement $f$ indirectly, despite preferences being private.
@@
Now, consider a new preference profile $R^{\prime}$ where, for every individual $i$, any outcome $\beta$ that was less preferred than $\alpha$ under $R_{i}$ remains less preferred under $R^{\prime}_{i}$. Observe that the strategy profile $s$ remains a Nash equilibrium of the mechanism $(S,g)$ under $R^{\prime}$ as well, because $\alpha$ is at least as preferred by each individual under $R^{\prime}$ as it was under $R$. Otherwise, if $s$ were no longer a Nash equilibrium, there would exist a profitable deviation for some individual under $R^{\prime}$, which would also represent a profitable deviation under $R$—a contradiction.
@@
The introduction of the NRMP centralized the matching process, enabling students and hospitals to submit rank-order lists of their preferences. The system used an algorithm similar to the institution-proposing Deferred Acceptance (DA) algorithm, independently discovered more than a decade later by Gale and Shapley (1962). [^19] The equivalence between the NRMP’s matching algorithm and the institution-proposing DA algorithm was identified in Roth (1984), decades after both were invented. Since Gale and Shapley formulated their celebrated algorithm without knowledge of the NRMP’s algorithm, I regard their equivalence as external validation of their invention. It aimed to produce stable outcomes, ensuring that no student and hospital would prefer each other over their assigned matches. [^20] See Section 6.4 for an in-depth discussion of stability and the DA algorithm. This innovation brought order and efficiency to the residency placement process for several decades.
@@
In Roth (2002), Roth clearly articulated the objective of his commission:
@@ 3.2.3 Pasteur’s Quadrant: Vickey Auction and Revenue Equivalence Theorem
“ [$\cdots$] it is of interest to inquire whether there is not some sealed-bid procedure that would be logically isomorphic to the progressive [English] auction. It is easily shown that the required procedure is to ask for bids on the understanding that the award will be made to the highest bidder, but on the basis of the price set by the second highest bidder.”
@@
Again, this strategic equivalence becomes evident when we interpret each buyer’s bid in the second-price auction as the maximum amount they are willing to bid in the English auction. With this breakthrough, Vickrey devised the second-price (sealed-bid) auction as a minimalist reform of the first-price auction, creating a static and strategy-proof auction format. [^24] Here, the strategy-proofness of the second-price auction applies to the bidders. However, a significant drawback of the second-price auction is its vulnerability to manipulation by the seller (Rothkopf et al., 1990; Akbarpour and Li, 2020). For instance, the seller could introduce fake bids just below the highest bid to artificially raise the winning price. Other limitations of the second-price auction include its vulnerability to collusion by a coalition of bidders (Ausubel and Milgrom, 2005).
@@
Though Vickrey was the first to introduce the second-price auction to scholarly literature, he was not the first to invent it (Moldovanu and Tietzel, 1998; Lucking-Reiley, 2000). In fact, the second-price auction had been invented and reinvented multiple times before Vickrey, likely because it represents a minimalist reform of both the first-price auction and the English auction. [^25] Since Vickrey formulated the second-price auction independently, without knowledge of its earlier inventions, and was the first to discuss it in academic literature, I regard these earlier instances as external validation of Vickrey’s invention.
@@
“I am inclined to offer Mr. Vieweg from Berlin an epic poem, Hermann and Dorothea, which will have approximately 2000 hexameters. [$\cdots$] Concerning the royalty we will proceed as follows: I will hand over to Mr. Counsel Böttiger a sealed note which contains my demand, and I wait for what Mr. Vieweg will suggest to offer for my work. If his offer is lower than my demand, then I take my note back, unopened, and the negotiation is broken. If, however, his offer is higher, then I will not ask for more than what is written in the note to be opened by Mr. Böttiger.”
@@
3.3 Rethinking Innovation Frameworks: Discovery-Invention Cycle
Minimalist market design integrates use-inspired basic research with policy initiatives, positioning it firmly within Pasteur’s Quadrant. Yet, it aligns even more closely with another framework that also challenges Bush’s linear model: the Discovery-Invention Cycle (DIC) model by Narayanamurti and Odumosu (2016). This model shapes the structure of this monograph, tracing the evolution of innovations across diverse applications in Sections 5 through 12 and highlighting the “discovery-invention cycles” that connect them.
@@ 3.3 Rethinking Innovation Frameworks: Discovery-Invention Cycle
According to this distinction, much of the work in commissioned market design projects, performed as a scheduled activity leading to the adopted design and “often required to be fast,” according to Roth (2002), is considered “development” rather than “research,” unlike the work performed after its completion as an unscheduled activity. However, in the spirit of Peter Medawar’s concerns discussed in Section 1.2, this research reflects post hoc research progress after the practical or policy impact has already occurred. Thus, it is less clear how the adapted design in commissioned settings may be received in cases where the design economist is an outsider advocate, since it is not stress-tested in such situations. Indeed, this has been a key reason why I sought an alternative approach to market design as a junior economist.
@@
To support their DIC model, Narayanamurti and Odumosu (2016) trace the evolution of information and communication technologies. They highlight Nobel Prize-winning achievements, such as the 1956 discovery of the transistor effect, which demonstrate several discovery-invention cycles over more than four decades. These cases show how, in some instances, scientific discoveries lead to inventions, while in others, inventions drive fundamental scientific breakthroughs. They write:
@@
Similarly, this monograph will illustrate how minimalist market design can lead to policy breakthroughs driven by use-inspired basic theory in some cases, while in others, policy interactions drive advancements in theory. By organizing Sections 5 through 12 to showcase a series of discovery-invention cycles in the theory and practice of matching markets, I will demonstrate how policy innovations can spur further theoretical developments and vice versa, akin to breakthroughs in information and communication technologies.
@@
“[$\cdots$] I am also not suggesting that social scientists do not already devote themselves to solving practical problems: many do, especially in policy-relevant areas like education, health care, poverty and government. Rather, what I am suggesting is that social scientists can profitably view the solution of practical problems as a mechanism for improving the coherency of social science itself.”
@@
The organization of Sections 5 through 12 demonstrates how theory and practice support each other in discovery-invention cycles, directly aligning with Watts’ theses. Narayanamurti and Odumosu suggest that “research should be evaluated not only by its initial motivations, but also by its ability to catalyze other research and innovations.” Through this monograph’s structure, the role of minimalist market design in catalyzing a series of innovations becomes apparent, including those that culminated during the COVID-19 pandemic when these innovations were most urgently needed.
Drawing inspiration from industrial laboratories like Bell Labs, Narayanamurti and Odumosu argue that Nobel Prize-winning industrial research undermines Bush’s notions of applied and basic research. They highlight that in the industrial research environment, “the very best discovery-oriented research was carried out.” Emphasizing the importance of connections between research and practice, they assert that the most productive distinction was not between basic and applied research but between various unscheduled research activities and scheduled product development.
Consequently, inspired by the workings of Bell Labs and insights from the DIC model, they propose policy suggestions for building research institutions that excel in innovation. Along similar lines, one of the central theses of this monograph argues that minimalist market design can achieve the same goal for teams of design economists by advancing theory and informing policy.
4 Advice for Design Economists with Policy Aspirations
Having elaborated on the role of use-inspired theory in minimalist market design in the previous section, I will now offer guidance for design economists seeking to adopt this framework as a policy instrument.
@@ 4.1 Pursuit of Promising Applications
Which practical applications are promising candidates for minimalist market design? How can a design economist identify them? Here, I offer my perspective.
@@
    Are you commissioned by decision makers or other stakeholders seeking changes to the existing system, or is it merely an aspiration of an outsider?
@@
This framework is especially useful when the design economist acts as an external critic, intervening in an existing institution that, while flawed, does not require a complete overhaul, and addressing its limitations that go beyond the textbook considerations of economics. Its cautious approach is particularly valuable in areas outside the traditional domain of economists, where interventions risk being perceived as intrusive.
At times, finding promising applications feels like detective work. Institutions often begin with simple missions that have straightforward solutions. However, as these missions evolve, the solutions can become more complex, especially for decision makers who lack training in analytical methods. In response to changing missions, decision makers may resort to patchwork modifications of simpler solutions, which can introduce various issues into the system. By tracing this history, a design economist can not only identify a candidate institution for improvement but also uncover the root cause of its failures.
Let’s illustrate this strategy with a hypothetical scenario. [^30] Conceptually, this scenario closely parallels the evolution of the U.S. Army’s cadet-branch matching mechanism, discussed in Section 8, as well as a high school admissions policy implemented across China from the early 1990s to the mid-2010s, analyzed in Zhou and Wang (2024).
@@
Even if the system operator is highly skilled and finds the correct solution in this scenario—which, in my experience, would be rare—a subsequent mission change could make the solution even more challenging. For instance, suppose several elite schools decide to allocate their seats jointly based on these principles while also considering students’ preferences. If our system operator avoided errors with the initial change, they will surely struggle with this more complex version of the problem. By tracing the institution’s history, our design economist can act like a detective, identifying such market failures and uncovering their root causes.
@@ 4.2 A Roadmap for an Aspiring Market Designer
Step 1. Identify a Real Problem: This is a substantial undertaking, so following this framework makes sense only if your scholarly efforts have a practical counterpart. Can you articulate the problem and your insights to friends who are not academic economists? Can you envision crafting a news story based on these ideas? Naturally, the more significant the problem, the better.
Step 2. Understand the Problem Inside and Out: To effectively address a real-life problem, comprehensive understanding is crucial. If all goes well, you will eventually engage with authorities and various stakeholders to advocate for your practical ideas. In many cases, your efforts might be perceived as a “liability” if they expose the limitations of an existing institution. Interactions with authorities who may have designed the “flawed” institution are inevitable, and they may be inclined to discredit and dismiss your efforts. Therefore, as a minimum requirement, you must have an in-depth understanding of the issue to reasonably expect to convince stakeholders that you can be a valuable “asset.”
Success in this endeavor demands familiarity not only with literature related to the application but also with other relevant materials from the field. For instance, before engaging decision makers and experts in other fields for our various policy-oriented projects, our teams of design economists:
@@
Step 3. Identify Key Policy Objectives of Stakeholders: The interests of academic economists may not always align with the critical concerns in real-world settings. Conducting market design research without thoroughly understanding stakeholders’ primary objectives risks producing insights or instruments of limited practical value. If your primary goal is to generate research that informs policy, it is paramount to understand the mission of the institution in question.
@@
Do not assume that the objectives of policymakers or stakeholders need to resemble standard assumptions in neoclassical economics. At this phase, thinking like a mainstream economist may not always serve you well. Neoclassical analysis largely relies on a preference utilitarian framework, where the primary objective is the maximization of some aggregate measure of preference satisfaction. Among my successful endeavors in market design, kidney exchange (see Section 7) and liver exchange (see Section 12) stand out as the only ones where preference utilitarianism effectively captures a key stakeholder objective. [^32] Even this statement comes with a caveat. Given the absence of monetary transfers in kidney and liver exchanges, the key desideratum in these applications is not utilitarianism but rather its ordinal variant, the Pareto principle (Broome, 2001). In these cases, a naturally fitting and objectively significant welfare measure revolves around the number of lives saved through donor exchanges. Consequently, the welfare gains presented in Roth et al. (2004, 2005b, 2007); Ergin et al. (2020) received a favorable response from members of the transplantation community.
In contrast, in my work on school choice, it was predominantly considerations tied to incentive compatibility that propelled a series of reforms (see Section 6). It was largely equity considerations that were the driving force for bioethicists and emergency healthcare experts endorsing and adopting our policy recommendations for addressing the pandemic allocation of scarce medical resources during COVID-19 (see Section 11). Regarding India’s affirmative action policies, recent reforms aligned with those advocated in Sönmez and Yenmez (2022) were exclusively steered by equity considerations (see Section 10). Lastly, both equity and incentive compatibility considerations played central roles in the U.S. Army’s reform of its branching process at West Point and ROTC (see Section 8).
@@
Step 4. Formulate a Realistic and Analytically Tractable Model: Understanding the problem thoroughly and identifying key policy objectives are just the beginning. Now, you need a model that is both realistic and analytically manageable. Determine which elements of the problem are essential and which are superfluous. Make your model as lean as possible without omitting any key considerations. Often, the structure of the existing institution and the policy goals you identified will be instrumental in shaping your model.
Step 5. Evaluate the Suitability of the Original Institution for Key Objectives: Your best bet for capturing the attention of various stakeholders is to demonstrate that the existing institution falls short in meeting at least one of their key objectives. Clearly, if the existing institution already aligns well with the key objectives of stakeholders, a reform is not necessary. Nevertheless, the overarching framework remains valuable for studying the existing system. After all, researchers in market design are not solely focused on achieving direct policy impact.
Step 6. Identify Root Causes of Any Failures: Even when an existing institution falls short of meeting some of the most critical stakeholder objectives, convincing them to embrace a costly change is a formidable task. Stakeholders are particularly unlikely to support a reform that entails a complete overhaul of the entire institution. Beyond the heightened difficulty and cost associated with more extensive reform, many elements within an institution may serve additional purposes you are unaware of. Therefore, it becomes imperative to enhance the institution while preserving its core structure. To achieve this, diagnosing the “root causes” of the failures becomes crucial. This stage is also valuable for optimizing your model. For instance, you can establish the primitives of the model based on which elements of the existing institution can be retained and which ones need to be modified or eliminated.
Step 7. Tackle the Root Causes of Failures to Develop an Alternative System Aligned with Key Objectives: This step embodies a defining aspect of “minimalist” market design, lending this paradigm its name.
@@
In rare cases, there will be a unique design that satisfies the desiderata (see Theorem 10 in Section 8.8.2 and Theorem 12 in Section 10.5). This is considered the gold standard in axiomatic methodology, provided that all axioms correspond to actual desiderata. Naturally, the desired reform is most evident in those cases. [^34] The value of uniqueness is less clear in practical applications than in scholarly inquiry. While a uniqueness result is valuable in research, partly because it avoids selection concerns and partly due to its elegance, the added flexibility of multiple solutions can be beneficial in practical applications.
Step 8. Solidify the Practical Value and Outreach of Your Proposal: If you have come this far, your research is of potential value to stakeholders. Before engaging them to explore their receptiveness to your policy analysis, it is advisable to support your conceptual and theoretical ideas with experimental, computational, or empirical methods. Additionally, gathering as much field or anecdotal evidence as possible can demonstrate the relevance of your research.

### v4

@@ 3.2.1 Bohr’s Quadrant: Nash Implementation
Maskin examines the following abstract setting: There is a set of individuals II and a set of outcomes $A$, where each individual $i\in I$ has a weak preference relation $R_{i}$ over $A$. For each individual $i\in I$, let $\mathcal{R}_{i}$ represent the set of possible preferences. Let $R=(R_{i})_{i\in I}$ denote a generic preference profile, and let $\mathcal{R}$ denote the set of all possible preference profiles. Importantly, preferences are private information.
A central planner aims to select an outcome based on individuals’ true preferences, formalized as a social choice function $f:\mathcal{R}\to A$ that maps each preference profile $R\in\mathcal{R}$ to an outcome $\alpha\in A$. [^20] Maskin’s theory is more general, allowing for social choice correspondences that map each preference profile to a set of outcomes.
@@
To overcome this, the planner designs a mechanism $(S,g)$ à la Hurwicz, where $S_{i}$ represents the strategy space for each individual $i\in I$, and $g:S\to A$ is the outcome function selecting an outcome $\alpha\in A$ for each strategy profile $s=(s_{i})_{i\in I}\in S=\prod_{i\in I}S_{i}$. The objective is to design a mechanism such that, for any preference profile $R$, the Nash equilibrium outcomes of the induced game coincide with the desired outcome $f(R)$. This allows the planner to implement $f$ indirectly, despite preferences being private.
@@
Now, consider a new preference profile $R^{\prime}$ where, for every individual $i$, any outcome $\beta$ that was less preferred than $\alpha$ under $R_{i}$ remains less preferred under $R^{\prime}_{i}$. Observe that the strategy profile $s$ remains a Nash equilibrium of the mechanism $(S,g)$ under $R^{\prime}$ as well, because $\alpha$ is at least as preferred by each individual under $R^{\prime}$ as it was under $R$. Otherwise, if $s$ were no longer a Nash equilibrium, there would exist a profitable deviation for some individual under $R^{\prime}$, which would also represent a profitable deviation under RR—a contradiction.
@@
The introduction of the NRMP centralized the matching process, enabling students and hospitals to submit rank-order lists of their preferences. The system used an algorithm similar to the institution-proposing deferred acceptance (DA) algorithm, independently discovered more than a decade later by Gale and Shapley (1962). [^23] The equivalence between the NRMP’s matching algorithm and the institution-proposing DA algorithm was identified in Roth (1984), decades after both were invented. Since Gale and Shapley formulated their celebrated algorithm without knowledge of the NRMP’s algorithm, I regard their equivalence as external validity for their invention. It aimed to produce stable outcomes, ensuring that no student and hospital would prefer each other over their assigned matches. [^24] See Section 6.4 for an in-depth discussion of stability and the DA algorithm. This innovation brought order and efficiency to the residency placement process for several decades.
@@
Roth (2002) clearly articulated the objective of his commission:
@@ 3.2.3 Pasteur’s Quadrant: Vickey Auction and Revenue Equivalence Theorem
“ $\cdots$] it is of interest to inquire whether there is not some sealed-bid procedure that would be logically isomorphic to the progressive [English] auction. It is easily shown that the required procedure is to ask for bids on the understanding that the award will be made to the highest bidder, but on the basis of the price set by the second highest bidder.”
@@
Again, this strategic equivalence becomes evident when we interpret each buyer’s bid in the second-price auction as the maximum amount they are willing to bid in the English auction. With this breakthrough, Vickrey devised the second-price (sealed-bid) auction as a minimalist reform of the first-price auction, creating a static and strategy-proof auction format. [^28] Here, the strategy-proofness of the second-price auction applies to the bidders. However, a significant drawback of the second-price auction is its vulnerability to manipulation by the seller (Rothkopf et al., 1990; Akbarpour and Li, 2020). For instance, the seller could introduce fake bids just below the highest bid to artificially raise the winning price. Other limitations of the second-price auction include its vulnerability to collusion by a coalition of bidders (Robinson, 1985; Klemperer, 2002; Ausubel and Milgrom, 2005).
@@
Vickrey’s second-price auction (and close variants) has seen real-world use across several settings, but outcomes have been mixed—in part because the format can be vulnerable to collusion and other strategic behavior in multi-item environments (Robinson, 1985; Klemperer, 2002; Ausubel and Milgrom, 2005). New Zealand’s brief experience with the second-price rule for allocating radio-spectrum “management rights” in 1989–1990 is widely cited as the first large-scale governmental deployment of a pure Vickrey format (McMillan, 1994; Crandall, 1998). [^29] After several highly publicized anomalies (e.g., a top bid of NZ\$100,000 yet payment of only NZ\$6 because there was no reserve price and the second-highest bid was NZ\$6), and thin bidding in some lots, New Zealand switched to first-price auction in 1991(McMillan, 1994; Ministry of Economic Development, 2005). Following international practice, New Zealand later moved to simultaneous ascending (multiple-round) auctions in 1996 (Ministry of Economic Development, 2005).
@@
Though Vickrey was the first to introduce the second-price auction to scholarly literature, he was not the first to invent it (Moldovanu and Tietzel, 1998; Lucking-Reiley, 2000). In fact, the second-price auction had been invented and reinvented multiple times before Vickrey, likely because it represents a minimalist reform of both the first-price auction and the English auction. [^30] Since Vickrey formulated the second-price auction independently, without knowledge of its earlier inventions, and was the first to discuss it in academic literature, I regard these earlier instances as external validity for Vickrey’s invention.
@@
“I am inclined to offer Mr. Vieweg from Berlin an epic poem, Hermann and Dorothea, which will have approximately 2000 hexameters. $\cdots$] Concerning the royalty we will proceed as follows: I will hand over to Mr. Counsel Böttiger a sealed note which contains my demand, and I wait for what Mr. Vieweg will suggest to offer for my work. If his offer is lower than my demand, then I take my note back, unopened, and the negotiation is broken. If, however, his offer is higher, then I will not ask for more than what is written in the note to be opened by Mr. Böttiger.”
@@
3.3 Rethinking Innovation Frameworks: Discovery–Invention Cycle
Minimalist market design integrates use-inspired basic research with policy initiatives, positioning it firmly within Pasteur’s Quadrant. Yet, it aligns even more closely with another framework that also challenges Bush’s linear model: the Discovery–Invention Cycle (DIC) model by Narayanamurti and Odumosu (2016). This model shapes the structure of this monograph, tracing the evolution of innovations across diverse applications in Sections 5 through 12 and highlighting the “discovery–invention cycles” that connect them, as illustrated in Figures 36 and 37.
@@ 3.3 Rethinking Innovation Frameworks: Discovery–Invention Cycle
According to this distinction, much of the work in commissioned market design projects, performed as a scheduled activity leading to the adopted design and “often required to be fast,” according to Roth (2002), is considered “development” rather than “research,” unlike the work performed after its completion as an unscheduled activity. However, in the spirit of Peter Medawar’s concerns discussed in Section 1.2, this research reflects post hoc research progress after the practical or policy impact has already occurred. Thus, it is less clear how the adapted design in commissioned settings may be received in cases where the economic designer is an outsider advocate, since it is not stress-tested in such situations. Indeed, this has been a key reason why I sought an alternative approach to market design as a junior economist.
@@
To support their DIC model, Narayanamurti and Odumosu (2016) trace the evolution of information and communication technologies. They highlight Nobel Prize-winning achievements, such as the 1956 discovery of the transistor effect, which demonstrate several discovery–invention cycles over more than four decades. These cases show how, in some instances, scientific discoveries lead to inventions, while in others, inventions drive fundamental scientific breakthroughs. They write:
@@
This monograph similarly illustrates how minimalist market design can lead to policy breakthroughs driven by use-inspired basic theory in some cases, while in others, policy interactions stimulate new theoretical advances. By organizing Sections 5 through 12 to showcase a series of discovery–invention cycles in the theory and practice of matching markets, as illustrated in Figures 36 and 37, I demonstrate how policy innovations can spur further theoretical developments—and vice versa—analogous to breakthroughs in information and communication technologies. [^34] The idea of a reciprocal relationship between theoretical discovery and practical invention in market design is not original to this monograph. A similar perspective—developed independently of the DIC model—appears in Pycia (2019), who emphasizes the two-way relationship between technological innovation and practical market design: new technologies enable novel designs, while new designs in turn catalyze the emergence of further technologies. Drawing on examples from medical, energy, automotive, computing, and data technologies, Pycia highlights how these interactions foster innovation in the theory and practice of market design. The present monograph builds on this broader reciprocal perspective within minimalist market design, linking use-inspired basic theory with policy reform across multiple domains.
@@
$\cdots$] I am also not suggesting that social scientists do not already devote themselves to solving practical problems: many do, especially in policy-relevant areas like education, health care, poverty and government. Rather, what I am suggesting is that social scientists can profitably view the solution of practical problems as a mechanism for improving the coherency of social science itself.”
@@
The organization of Sections 5 through 12 demonstrates how theory and practice support each other in discovery–invention cycles, directly aligning with Watts’ theses. Narayanamurti and Odumosu suggest that “research should be evaluated not only by its initial motivations, but also by its ability to catalyze other research and innovations.” Through this monograph’s structure, the role of minimalist market design in catalyzing a series of innovations becomes apparent, including those that culminated during the COVID-19 pandemic when these innovations were most urgently needed.
Drawing inspiration from industrial laboratories like Bell Labs, Narayanamurti and Odumosu argue that Nobel Prize-winning industrial research undermines Bush’s notions of applied and basic research. They highlight that in the industrial research environment, “the very best discovery–oriented research was carried out.” Emphasizing the importance of connections between research and practice, they assert that the most productive distinction was not between basic and applied research but between various unscheduled research activities and scheduled product development.
Consequently, inspired by the workings of Bell Labs and insights from the DIC model, they propose policy suggestions for building research institutions that excel in innovation. Along similar lines, one of the central theses of this monograph argues that minimalist market design can achieve the same goal for teams of economic designers by advancing theory and informing policy.
4 Advice for Economic Designers with Policy Aspirations
Having elaborated on the role of use-inspired theory in minimalist market design in the previous section, I will now offer guidance for economic designers seeking to adopt this framework as a policy instrument.
@@ 4.1 Pursuit of Promising Applications
Which practical applications are promising candidates for minimalist market design? How can an economic designer identify them? Here, I offer my perspective.
@@
    Are you commissioned by decision-makers or other stakeholders seeking changes to the existing system, or is it merely an aspiration of an outsider?
@@
This framework is especially useful when the economic designer acts as an external critic, intervening in an existing institution that, while flawed, does not require a complete overhaul, and addressing its limitations that go beyond the textbook considerations of economics. Its cautious approach is particularly valuable in areas outside the traditional domain of economists, where interventions risk being perceived as intrusive.
At times, finding promising applications feels like detective work. Institutions often begin with simple missions that have straightforward solutions. However, as these missions evolve, the solutions can become more complex, especially for decision-makers who lack training in analytical methods. In response to changing missions, decision-makers may resort to patchwork modifications of simpler solutions, which can introduce various issues into the system. By tracing this history, an economic designer can not only identify a candidate institution for improvement but also uncover the root cause of its failures.
Let’s illustrate this strategy with a hypothetical scenario. [^36] Conceptually, this scenario closely parallels the evolution of the U.S. Army’s cadet–branch matching mechanism, discussed in Section 8, as well as a high school admissions policy implemented across China from the early 1990s to the mid-2010s, analyzed in Zhou and Wang (2024).
@@
Even if the system operator is highly skilled and finds the correct solution in this scenario—which, in my experience, would be rare—a subsequent mission change could make the solution even more challenging. For instance, suppose several elite schools decide to allocate their seats jointly based on these principles while also considering students’ preferences. If our system operator avoided errors with the initial change, they will surely struggle with this more complex version of the problem. By tracing the institution’s history, our economic designer can act like a detective, identifying such market failures and uncovering their root causes.
@@ 4.2 A Roadmap for an Aspiring Market Designer
Step 1. Identify a Real Problem.
This is a substantial undertaking, so following this framework makes sense only if your scholarly efforts have a practical counterpart. Can you articulate the problem and your insights to friends who are not academic economists? Can you envision crafting a news story based on these ideas? Naturally, the more significant the problem, the better.
Step 2. Understand the Problem Inside and Out.
To effectively address a real-life problem, comprehensive understanding is crucial. If all goes well, you will eventually engage with authorities and various stakeholders to advocate for your practical ideas. In many cases, your efforts might be perceived as a “liability” if they expose the limitations of an existing institution. Interactions with authorities who may have designed the “flawed” institution are inevitable, and they may be inclined to discredit and dismiss your efforts. Therefore, as a minimum requirement, you must have an in-depth understanding of the issue to reasonably expect to convince stakeholders that you can be a valuable “asset.”
Success in this endeavor demands familiarity not only with literature related to the application but also with other relevant materials from the field. For instance, before engaging decision-makers and experts in other fields for our various policy-oriented projects, our teams of economic designers:
@@
Step 3. Identify Key Policy Objectives of Stakeholders.
The interests of academic economists may not always align with the critical concerns in real-world settings. Conducting market design research without thoroughly understanding stakeholders’ primary objectives risks producing insights or instruments of limited practical value. If your primary goal is to generate research that informs policy, it is paramount to understand the mission of the institution in question.
@@
Do not assume that the objectives of policymakers or stakeholders need to resemble standard assumptions in neoclassical economics. At this phase, thinking like a mainstream economist may not always serve you well. Neoclassical analysis largely relies on a preference utilitarian framework, where the primary objective is the maximization of some aggregate measure of preference satisfaction. Among my successful endeavors in market design, kidney exchange (see Section 7) and liver exchange (see Section 12) stand out as the only ones where preference utilitarianism effectively captures a key stakeholder objective. In these cases, a naturally fitting and objectively significant welfare measure revolves around the number of lives saved through donor exchanges. Consequently, the welfare gains presented in Roth et al. (2004, 2005b, 2007); Ergin et al. (2020) received a favorable response from members of the transplantation community.
In contrast, in my work on school choice, it was predominantly considerations tied to incentive compatibility that propelled a series of reforms (see Section 6). It was largely equity considerations that were the driving force for bioethicists and emergency healthcare experts endorsing and adopting our policy recommendations for addressing the pandemic allocation of scarce medical resources during COVID-19 (see Section 11). Regarding India’s affirmative action policies, recent reforms aligned with those advocated in Sönmez and Yenmez (2022) were exclusively steered by equity considerations (see Section 10). Lastly, both equity and incentive compatibility considerations played central roles in the U.S. Army’s reform of its branching process at West Point and ROTC (see Section 8). [^38] Kominers et al. (2017) similarly highlight that equity considerations play a fundamental role across many market design settings.
@@
Step 4. Formulate a Realistic and Analytically Tractable Model.
Understanding the problem thoroughly and identifying key policy objectives are just the beginning. Now, you need a model that is both realistic and analytically manageable. Determine which elements of the problem are essential and which are superfluous. Make your model as lean as possible without omitting any key considerations. Often, the structure of the existing institution and the policy goals you identified will be instrumental in shaping your model.
Step 5. Evaluate the Suitability of the Original Institution for Key Objectives.
Your best bet for capturing the attention of various stakeholders is to demonstrate that the existing institution falls short in meeting at least one of their key objectives. Clearly, if the existing institution already aligns well with the key objectives of stakeholders, a reform is not necessary. Nevertheless, the overarching framework remains valuable for studying the existing system. After all, researchers in market design are not solely focused on achieving direct policy impact.
Step 6. Identify Root Causes of Any Failures.
Even when an existing institution falls short of meeting some of the most critical stakeholder objectives, convincing them to embrace a costly change is a formidable task. Stakeholders are particularly unlikely to support a reform that entails a complete overhaul of the entire institution. Beyond the heightened difficulty and cost associated with more extensive reform, many elements within an institution may serve additional purposes you are unaware of. Therefore, it becomes imperative to enhance the institution while preserving its core structure. To achieve this, diagnosing the “root causes” of the failures becomes crucial. This stage is also valuable for optimizing your model. For instance, you can establish the primitives of the model based on which elements of the existing institution can be retained and which ones need to be modified or eliminated.
Step 7. Tackle the Root Causes of Failures to Develop an Alternative System Aligned with Key Objectives.
This step embodies a defining aspect of “minimalist” market design, lending this paradigm its name.
@@
In rare cases, there will be a unique design that satisfies the desiderata (see, for example, Theorem 10 in Section 8.8.2 and Theorem 12 in Section 10.5). This is considered the gold standard in axiomatic methodology, provided that all axioms correspond to actual desiderata. Naturally, the desired reform is most evident in those cases. [^40] The value of uniqueness is less clear in practical applications than in scholarly inquiry. While a uniqueness result is valuable in research, partly because it avoids selection concerns and partly due to its elegance, the added flexibility of multiple solutions can be beneficial in practical applications.
Step 8. Solidify the Practical Value and Outreach of Your Proposal.
If you have come this far, your research is of potential value to stakeholders. Before engaging them to explore their receptiveness to your policy analysis, it is advisable to support your conceptual and theoretical ideas with experimental, computational, or empirical methods. Additionally, gathering as much field or anecdotal evidence as possible can demonstrate the relevance of your research.